# SciReplicate-Bench: Benchmarking LLMs in Agent-driven Algorithmic Reproduction from Research Papers

### v1

@@ Abstract
This study evaluates large language models (LLMs) in generating code from algorithm descriptions from recent NLP papers. The task requires two key competencies: (1) algorithm comprehension: synthesizing information from papers and academic literature to understand implementation logic, and (2) coding expertise: identifying dependencies and correctly implementing necessary APIs. To facilitate rigorous evaluation, we introduce SciReplicate-Bench, a benchmark of 100 tasks from 36 NLP papers published in 2024, featuring detailed annotations and comprehensive test cases. Building on SciReplicate-Bench, we propose Sci-Reproducer, a multi-agent framework consisting of a Paper Agent that interprets algorithmic concepts from literature and a Code Agent that retrieves dependencies from repositories and implement solutions. To assess algorithm understanding, we introduce reasoning graph accuracy, which quantifies similarity between generated and reference reasoning graphs derived from code comments and structure. For evaluating implementation quality, we employ execution accuracy, CodeBLEU, and repository dependency/API recall metrics. In our experiments, we evaluate various powerful Non-Reasoning LLMs and Reasoning LLMs as foundational models. The best-performing LLM using Sci-Reproducer achieves only 39% execution accuracy, highlighting the benchmark’s difficulty. Our analysis identifies missing or inconsistent algorithm descriptions as key barriers to successful reproduction. We will open-source our benchmark, and code at https://github.com/xyzCS/SciReplicate-Bench.
@@ 1 Introduction
Among the aforementioned advancements in research acceleration, the ability of LLMs to correctly generate code for validating real-world scientific ideas is particularly noteworthy. Computational validation is crucial across many fields, yet researchers often face barriers due to limited coding expertise or inaccessible implementations. By converting scientific algorithm descriptions into executable code, LLMs could enhance reproducibility and accelerate scientific discovery. However, despite progress in LLM-based code generation, a significant gap remains in generating code directly from scholarly papers. Firstly, algorithm comprehension from scientific papers is challenging. Research papers are characterized by their brevity, methodological rigor, and extensive citations, with critical details about algorithms often dispersed across multiple sections of the paper. Understanding these algorithms requires synthesizing information from internal references and external scholarly works. Secondly, code repositories typically comprise multiple interdependent files and directories. To implement an algorithm, LLMs must comprehensively examine file dependencies, identify reusable components, and correctly handle both internal dependencies and external APIs.
Despite the importance of automatic scientific ideas verification, there exists no dataset specifically designed to evaluate the ability of LLMs to reproduce real-world algorithms proposed in peer-reviewed publications. As shown in Table 1, there are several machine learning software engineering benchmarks primarily focused on evaluating algorithmic design or straightforward implementations, which are significantly less complex than the methods typically described in academic research papers. For example, MLE-BENCH (Liu et al., 2023) and MLAgentBench (Huang et al., 2023) utilize Kaggle competitions, where LLMs must develop and implement solutions based on provided task specifications. ML-BENCH Chan et al. (2024) uses Machine Learning (ML) GitHub repositories to assess LLMs’ text-to-code capabilities and test autonomous agents in task execution.
@@
Therefore, we manually developed SciReplicate-Bench, the first benchmark specifically designed to evaluate LLMs’ capabilities in code generation for reproducing research findings from academic papers. It consists of 100 code reproduction tasks derived from 36 papers published in top NLP conferences in 2024. This recent publication window was deliberately chosen to minimize the risk of data leakage. An overview of the task is illustrated in Figure 1, with a concrete example provided in Figure A2 in Appendix E. The task consists of two main steps: 1. Algorithm understanding. LLMs must extract essential information from the paper, such as workflow details, algorithm descriptions, and hyperparameter values. 2. Code implementation. Based on the extracted information, LLMs are required to implement a function or method within a provided repository, based on the LaTeX representation of the algorithm from the paper.
To rigorously assess LLM performance on this benchmark, we evaluate two dimensions corresponding to aforementioned two steps: algorithm comprehension correctness and code correctness. To evaluate algorithm comprehension, we introduce a reasoning graph to represent the reasoning logic behind algorithm reproduction. Each node in the graph represents a code comment, which reflects a single reasoning step and is aligned with a specific segment of code. Edges between nodes are defined based on data flow relationships across different code segments. We compute the similarity between the generated reasoning graph and a reference graph to derive the reasoning graph accuracy. To evaluate code correctness, we employ established metrics including execution accuracy (Rajkumar et al., 2022; Xiang et al., 2023), CodeBLEU (Ren et al., 2020), and recall of intra/cross-file dependencies and APIs.
@@
  • Benchmarks: we present SciReplicate-Bench, The first benchmark aims to evaluate LLMs’ ability in reproducing the algorithms proposed in real-world NLP publications.
  • Metric: we propose a reasoning graph accuracy for evaluating the correctness of implicit reasoning process behind code generation, serving as an indicator of an LLM’s algorithm comprehension capability.
  • Approach: we introduce Sci-Reproducer, a dual-agent framework that combines a paper agent for extracting and understanding algorithmic logic from research papers with a code agent that retrieves necessary dependencies from code repositories. Their collaborative enables comprehensive paper reproduction by combining deep algorithmic understanding with practical implementation capabilities.
  • Observations: we employed various state-of-the-art LLMs on this benchmark. The experimental results show that (i) even the most advanced models still find this task very challenging, with execution accuracy rates below 40% for all models. (ii) reasoning models often rely excessively on internal reasoning instead of utilizing pre-defined actions to extract relevant information from literature and code contexts. This behavior, referred to as the ‘Overthinking Problem’ (Cuadron et al., 2025; Sui et al., 2025), can lead to suboptimal performance in agent-based systems. (iii) LLMs can comprehend algorithm logic but face challenges with actual implementation. (iv) many failures stem from the missing or inconsistent algorithm descriptions, and the incorporation of the Paper Agent effectively mitigates issues related to incomplete information.
@@ 2 SciReplicate-Bench
SciReplicate-Bench is a benchmark consisting of 100 tasks curated from 36 NLP publications, constructed using their open-source code repositories and corresponding LaTeX sources. The task categories are detailed in Appendix A1. The benchmark focuses on repository-level code generation, where each task is centered around implementing a specific function or class method. As illustrated in Figure A3 in Appendix E, each task comprises nine components, which can be categorized into three groups corresponding to code generation, evaluation, and analysis, respectively.
@@
  2. Algorithm LaTex code: The LaTeX code description of the target algorithm, typically located within a subsection or paragraph of the target paper.
@@
  1. Implementation reference: the ground-truth implementation of the target algorithm, used for computing the CodeBLEU metric.
  2. Annotated reasoning graph: a structured graph capturing the reasoning process behind the reference implementation, used to evaluate reasoning graph accuracy.
  3. Dependency information: annotations covering intra-file and cross-file dependencies, as well as external APIs invoked in the reference code, used to calculate recall across all three dependency types.
  4. Verification suite: each task includes a dedicated Python environment with ten test cases and scripts to verify the correctness of the output produced by generated code.
@@
Based on SciReplicate-Bench, an LLM is given the algorithm LaTeX code, function signature, literature context, and repository context as input. The LLM is asked to output a function that implements the target algorithm.
@@
The benchmark construction process comprises four key steps: paper selection, Python environment setup, documentation, and verification suite preparation. To mitigate the risk of data leakage, we selected papers published in 2024 that provide publicly available code repositories. During the annotation process, each repository was refactored to isolate the core algorithm into a standalone function, and all sources of randomness were removed to ensure reproducibility and prevent leakage. On average, annotating each paper requires approximately 12 hours. Details of the annotation process are provided in Appendix B.
@@ 2.1.1 Evaluating Algorithm Comprehension
  2. Edge matching: for each reference edge $e_{r}=\langle n_{r}^{i},n_{r}^{j}\rangle$, if both endpoints are matched, we apply BFS to verify whether a corresponding edge exists in $G_{g}$.
@@ 3 Sci-Reproducer
To address this task, we introduce Sci-Reproducer, a dual-agent framework designed for scientific paper methodology replication. As illustrated in Figure 1, Sci-Reproducer  comprises a Paper Agent and a Code Agent that collaboratively work to replicate algorithms described in a given paper. The predefined actions employed by the agents are summarized in Table 2, with implementation details provided in Appendix D.
@@ 3.1 Paper Agent
Due to the input length limitations of LLMs, it is infeasible to input entire paper along with their associated literature. Consequently, the Paper Agent must selectively extract pertinent information, following a strategy akin to Retrieval Augmented Generation (RAG) (Wang et al., 2024; Sarthi et al., 2024). The Paper Agent incrementally builds an understanding of the target algorithm by executing predefined actions to query the literature context. To facilitate this process, we adopt ReAct (Yao et al., 2022) as the agent strategy, which enables seamless integration of action execution with intermediate reasoning steps.
After the Paper Agent concludes that all necessary information has been collected, it generates a comprehensive report comprising key findings that fill in the missing components of the target algorithm’s LaTeX source. An example of the report is shown in Figure A8. This report subsequently serves as a crucial input for the Code Agent. The prompt used to guide the Paper Agent is provided in Figure A6.
@@ 3.2 Code Agent
The Code Agent integrates the target algorithm’s LaTeX code with the Paper Agent’s report to comprehensively understand the algorithm. It leverages actions to search the code repository for necessary dependencies that aid implementation. The agent can also browse websites for additional information and use a compiler to test and iteratively debug the code, ensuring proper execution by identifying and fixing syntax errors. The prompt for the Code Agent is provided in Figure A7.
@@ 4 Experiments
We evaluate Sci-Reproducer on the SciReplicate-Bench  benchmark using 7 advanced LLMs, including five non-reasoning LLMs: GPT-4o-mini (4o mini, 2024), GPT-4o (GPT-4o, 2024), Claude-Sonnet-3.7 (Claude-Sonnet-3.7, 2025), Gemini-2.0-Flash (Gemini-2.0-Flash, 2024), and Deepseek-V3 (DeepSeek-AI et al., 2024), and different versions of the reasoning models O3-mini (o3 mini, 2024), i.e., three different levels of reasoning intensity. For the reasoning graph accuracy metric, node matching is performed using GPT-4o, which may introduce some randomness. To reduce this variability, we set the temperature to 0 and top-p to 1, ensuring deterministic generation. The calculation is repeated three times, and we report the average score as the final result.
@@
No Agent 0.030 0.238 0.708 0.012 0.000 0.217
Paper Agent 0.040 0.246 0.739 0.024 0.000 0.251
Code Agent 0.140 0.279 0.747 0.565 0.364 0.328
Sci-Reproducer 0.170 0.303 0.768 0.576 0.364 0.362
@@
No Agent 0.040 0.259 0.727 0.059 0.000 0.281
Paper Agent 0.020 0.263 0.732 0.023 0.000 0.298
Code Agent 0.260 0.325 0.803 0.682 0.576 0.421
Sci-Reproducer 0.270 0.329 0.808 0.688 0.636 0.417
@@
No Agent 0.070 0.282 0.739 0.094 0.091 0.362
Paper Agent 0.050 0.291 0.736 0.082 0.091 0.379
Code Agent 0.320 0.394 0.784 0.765 0.545 0.545
Sci-Reproducer 0.390 0.401 0.794 0.776 0.636 0.626
@@
Code Agent 0.220 0.323 0.725 0.553 0.212 0.426
Sci-Reproducer 0.250 0.346 0.758 0.588 0.333 0.455
@@
No Agent 0.030 0.260 0.747 0.012 0.061 0.272
Paper Agent 0.050 0.275 0.762 0.012 0.030 0.306
Code Agent 0.210 0.312 0.776 0.482 0.182 0.383
Sci-Reproducer 0.220 0.334 0.778 0.565 0.333 0.443
@@
No Agent 0.080 0.259 0.767 0.035 0.000 0.323
Paper Agent 0.050 0.262 0.768 0.035 0.000 0.315
Code Agent 0.150 0.278 0.815 0.306 0.000 0.348
Sci-Reproducer 0.180 0.280 0.806 0.376 0.121 0.328
@@
No Agent 0.040 0.263 0.747 0.035 0.000 0.336
Paper Agent 0.060 0.263 0.752 0.047 0.000 0.319
Code Agent 0.220 0.289 0.792 0.376 0.030 0.404
Sci-Reproducer 0.240 0.283 0.799 0.341 0.061 0.362
@@
No Agent 0.070 0.269 0.723 0.047 0.000 0.345
Paper Agent 0.070 0.267 0.774 0.035 0.000 0.366
Code Agent 0.160 0.277 0.797 0.165 0.152 0.374
Sci-Reproducer 0.160 0.283 0.810 0.294 0.091 0.357

### v2

@@ Abstract
[…]
@@ 1 Introduction
Among the aforementioned advancements in research acceleration, the ability of LLMs to correctly generate code for validating real-world scientific ideas is particularly noteworthy. Computational validation is crucial across many fields, yet researchers often face barriers due to limited coding expertise or inaccessible implementations. By converting scientific algorithm descriptions into executable code, LLMs could enhance reproducibility and accelerate scientific discovery. However, despite progress in LLM-based code generation, a significant gap remains in generating code directly from scholarly papers. First, algorithm comprehension from scientific papers is challenging. Research papers are characterized by their brevity, methodological rigor, and extensive citations, with critical details about algorithms often dispersed across multiple sections of the paper. Understanding these algorithms requires synthesizing information from internal references and external scholarly works. Second, code repositories typically consist of multiple interdependent files and directories. To implement an algorithm, LLMs must comprehensively examine file dependencies, identify reusable components, and correctly handle both internal dependencies and external APIs.
Despite the importance of automated scientific idea verification, there exists no dataset specifically designed to evaluate the ability of LLMs to reproduce real-world algorithms proposed in peer-reviewed publications. As shown in Table 1, there are several machine learning software engineering benchmarks primarily focused on evaluating algorithmic design or straightforward implementations, which are significantly less complex than the methods typically described in academic research papers. For example, MLE-BENCH (Liu et al., 2023) and MLAgentBench (Huang et al., 2023) utilize Kaggle competitions, where LLMs must develop and implement solutions based on provided task specifications. ML-BENCH (Chan et al., 2024) uses Machine Learning (ML) GitHub repositories to assess LLMs’ text-to-code capabilities and test autonomous agents in task execution.
@@
To address this gap, we developed SciReplicate-Bench, the first benchmark specifically designed to evaluate LLMs’ capabilities in code generation for reproducing research findings from academic papers. It consists of 100 code reproduction tasks derived from 36 papers published in leading conferences in 2024. This recent publication window was deliberately chosen to minimize the risk of data leakage. An overview of the task is illustrated in Figure 1, with a concrete example provided in Figure A2 in Appendix E. The task consists of two main steps: 1. Algorithm understanding. LLMs must extract essential information from the paper, such as workflow details, algorithm descriptions, and hyperparameter values. 2. Code implementation. LLMs then implement a function or method within a provided repository, using both the extracted information and the LaTeX representation of the algorithm from the paper. We introduce Sci-Reproducer, a dual-agent system that combines a Paper Agent and a Code Agent to handle these two steps collaboratively and implement code for the target algorithm.
To rigorously assess LLM performance on this benchmark, we evaluate two dimensions corresponding to the aforementioned two steps: algorithm comprehension correctness and code correctness. To evaluate algorithm comprehension, we introduce a reasoning graph to represent the reasoning logic behind algorithm reproduction. Each node in the graph represents a code comment, which reflects a single reasoning step and is aligned with a specific segment of code. Edges between nodes are defined based on data flow relationships across different code segments. We compute the similarity between the generated reasoning graph and a reference graph to derive the reasoning graph accuracy. To evaluate code correctness, we employ established metrics including execution accuracy (Rajkumar et al., 2022; Xiang et al., 2023), CodeBLEU (Ren et al., 2020), and recall of intra/cross-file dependencies and APIs.
@@
  • Benchmarks: SciReplicate-Bench, a benchmark of 100 algorithm reproduction tasks from recent NLP publications.
  • Metric: We propose a novel reasoning graph accuracy metric for evaluating algorithmic comprehension.
  • Approach: Sci-Reproducer, a dual-agent framework combining paper understanding and code implementation.
  • Insights: Comprehensive evaluation across state-of-the-art LLMs reveals four key findings: (i) the task remains highly challenging, with execution accuracy below 40% for all models; (ii) reasoning models exhibit “overthinking” behavior (Cuadron et al., 2025; Sui et al., 2025), over-relying on internal reasoning rather than utilizing available tools for information extraction; (iii) while LLMs demonstrate strong algorithmic comprehension, they struggle with practical implementation; and (iv) algorithm reproduction failures often stem from incomplete paper descriptions, which our Sci-Reproducer effectively addresses.
2 Related Work
Our work lies at the intersection of AI for automating scientific discovery and LLM-based code generation.
2.1 AI for Automating Scientific Discovery
The application of LLMs to accelerate scientific research has emerged as a rapidly growing field with diverse approaches. Several studies have demonstrated the potential for comprehensive research automation through end-to-end AI systems. Schmidgall et al. (2025); Lu et al. (2024) developed frameworks that integrate idea generation, experimental validation, and manuscript composition, with some AI-authored papers successfully passing workshop review processes (Yamada et al., 2025). Complementary research has focused on the creative aspects of scientific inquiry, with Wang et al. (2023); Ghafarollahi & Buehler (2024); O’Neill et al. (2025) investigating LLMs’ capacity for generating novel research hypotheses. Notably, recent evaluations by Gu et al. (2024); Kumar et al. (2024); Liu et al. (2025); Si et al. (2024) suggest that AI-generated research concepts may occasionally exceed human-generated ideas in terms of novelty and originality.
Within computational disciplines where implementation validation is essential, LLMs have shown promise in algorithm design and code development tasks. Our proposed SciReplicate-Bench addresses a previously underexplored area: the automated reproduction of algorithms directly from academic publications. This represents a unique challenge at the intersection of scientific literature comprehension and executable code synthesis. While recent parallel efforts such as PaperBench (Starace et al., 2025) and Paper2CodeBench (Seo et al., 2025) have tackled related problems by exploring full codebase reconstruction, these evaluation approaches rely substantially on manual assessment criteria and LLM-based correctness judgments, introducing potential inconsistencies and reliability concerns. Our approach prioritizes objective evaluation through execution accuracy, providing more rigorous validation than non-executable assessment methodologies (Wang et al., 2022; Chen et al., 2021).
The broader ecosystem of computational reproducibility research includes specialized frameworks such as MLGym (Nathani et al., 2025) for baseline improvement, and evaluation benchmarks developed by Siegel et al. (2024); Ren et al. (2023) that assess LLMs’ ability to reproduce published experimental results using existing codebases.
2.2 LLMs for Code Generation
Code generation has emerged as a prominent application of LLMs, with benchmarks ranging from basic programming tasks (Chen et al., 2021; Jain et al., 2024; Austin et al., 2021; Hendrycks et al., 2021; Liu et al., 2022) to realistic software engineering challenges like SWE-bench (Jimenez et al., 2023), which uses actual repository pull requests. However, these benchmarks primarily target general software engineering rather than scientific algorithm reproduction.
Recent efforts have developed machine learning-specific benchmarks (Liu et al., 2023; Huang et al., 2023; Chan et al., 2024), but these typically involve implementing algorithms proposed by the models themselves or solving relatively straightforward tasks. They lack the depth of algorithmic understanding and rigorous paper analysis required for reproducing algorithms from peer-reviewed publications.
Despite advances in tool-augmented code generation (Schick et al., 2023; Zhang et al., 2024a; a; 2023b), no existing system specifically addresses the unique challenge of translating academic papers into executable code. Our Sci-Reproducer framework demonstrates the ability to comprehend academic publications and convert abstract algorithm descriptions into functional implementations.
@@ 3 SciReplicate-Bench
SciReplicate-Bench is designed to evaluate LLMs’ ability to reproduce algorithms from academic papers, consisting of 100 tasks curated from 36 recent NLP publications with their corresponding open-source implementations. The task categories are detailed in Figure A1 in Appendix B. The benchmark focuses on repository-level code generation, where each task is centered around implementing a specific function or class method. As illustrated in Figure A3 in Appendix E, each task comprises nine components, which can be categorized into three groups corresponding to code generation, evaluation, and analysis, respectively.
@@
  2. Algorithm Description: The LaTeX code description of the target algorithm, typically located within a subsection or paragraph of the target paper.
@@
  1. Reference implementation: ground-truth code serving as the reference for CodeBLEU evaluation.
  2. Reasoning graph annotations: structured representations of the algorithmic logic and implementation flow, enabling assessment of reasoning graph accuracy.
  3. Dependency annotations: comprehensive documentation of internal dependencies, cross-file relationships, and external API usage for computing recall metrics across all dependency categories.
  4. Test environment: isolated Python execution environment containing validation cases and automated verification scripts for assessing implementation correctness.
@@
Based on SciReplicate-Bench, an LLM is given the algorithm description, function signature, literature context, and repository context as input. The LLM is asked to output a function that implements the target algorithm.
@@
The benchmark construction process comprises four key steps: paper selection, Python environment setup, documentation, and verification suite preparation. To mitigate the risk of data leakage, we selected papers published in 2024 that provide publicly available code repositories. During the annotation process, each repository was refactored to isolate the core algorithm into a standalone function, and all sources of randomness were removed to ensure reproducibility and prevent leakage. On average, annotating each paper requires approximately 12 hours. Details of the annotation process are provided in Appendix A1.
@@ 3.1.1 Evaluating Algorithm Comprehension
  2. Edge matching: for each reference edge $e_{r}=\langle n_{r}^{i},n_{r}^{j}\rangle$, if both endpoint nodes have corresponding matches in $G_{g}$, we apply Breadth-First Search(BFS) to verify whether a corresponding edge exists in $G_{g}$.
@@ 4 Sci-Reproducer
To address this task, we introduce Sci-Reproducer  [^4] A video demonstration showcasing Sci-Reproducer’s capabilities is provided at https://youtu.be/qcSIMgyehjE, a dual-agent framework designed for scientific paper algorithm replication. As illustrated in Figure 1, Sci-Reproducer  comprises a Paper Agent and a Code Agent that collaboratively work to replicate algorithms described in a given paper. The predefined actions employed by the agents are summarized in Table 2, with implementation details provided in Appendix C.
@@ 4.1 Paper Agent
Based on the provided algorithm description, the Paper Agent systematically retrieves contextual information from the literature context to support algorithmic understanding. Due to the input length limitations of LLMs, it is infeasible to input entire paper along with their associated literature. Consequently, the Paper Agent must selectively extract pertinent information, following a strategy akin to Retrieval Augmented Generation (RAG) (Wang et al., 2024b; Sarthi et al., 2024). The Paper Agent incrementally builds an understanding of the target algorithm by executing predefined actions to query the literature context. To facilitate this process, we adopt ReAct (Yao et al., 2022) as the agent strategy, which enables seamless integration of action execution with intermediate reasoning steps.
After the Paper Agent concludes that all necessary information has been collected, it generates a comprehensive literature report comprising key findings that fill in the missing components of the target algorithm’s LaTeX source. An example of the literature report is shown in Figure A8. This report subsequently serves as a crucial input for the Code Agent. The prompt used to guide the Paper Agent is provided in Figure A6.
@@ 4.2 Code Agent
Informed by the algorithm description, literature report, and code context, the Code Agent searches the repository to locate essential dependencies required for implementation. It can also browse websites for additional information and use a compiler to test and iteratively debug the code, ensuring proper execution by identifying and fixing syntax errors. The prompt for the Code Agent is provided in Figure A7.
@@ 5 Experiments
No Agent 0.030 0.238 0.694 0.012 0.000 0.217
Paper Agent 0.040 0.246 0.717 0.024 0.000 0.251
Code Agent 0.140 0.279 0.738 0.565 0.364 0.328
Sci-Reproducer 0.170 0.303 0.741 0.576 0.364 0.362
@@
No Agent 0.040 0.259 0.705 0.059 0.000 0.281
Paper Agent 0.020 0.263 0.724 0.023 0.000 0.298
Code Agent 0.260 0.325 0.748 0.682 0.576 0.421
Sci-Reproducer 0.270 0.329 0.751 0.688 0.636 0.417
@@
No Agent 0.070 0.282 0.725 0.094 0.091 0.362
Paper Agent 0.050 0.291 0.727 0.082 0.091 0.379
Code Agent 0.320 0.394 0.764 0.765 0.545 0.545
Sci-Reproducer 0.390 0.401 0.773 0.776 0.636 0.626
@@
Code Agent 0.220 0.323 0.708 0.553 0.212 0.426
Sci-Reproducer 0.250 0.346 0.727 0.588 0.333 0.455
@@
No Agent 0.030 0.260 0.712 0.012 0.061 0.272
Paper Agent 0.050 0.275 0.732 0.012 0.030 0.306
Code Agent 0.210 0.312 0.736 0.482 0.182 0.383
Sci-Reproducer 0.220 0.334 0.738 0.565 0.333 0.443
@@
No Agent 0.080 0.259 0.758 0.035 0.000 0.323
Paper Agent 0.050 0.262 0.729 0.035 0.000 0.315
Code Agent 0.150 0.278 0.803 0.306 0.000 0.348
Sci-Reproducer 0.180 0.280 0.771 0.376 0.121 0.328
@@
No Agent 0.040 0.263 0.729 0.035 0.000 0.336
Paper Agent 0.060 0.263 0.726 0.047 0.000 0.319
Code Agent 0.220 0.289 0.749 0.376 0.030 0.404
Sci-Reproducer 0.240 0.283 0.758 0.341 0.061 0.362
@@
No Agent 0.070 0.269 0.718 0.047 0.000 0.345
Paper Agent 0.070 0.267 0.751 0.035 0.000 0.366
Code Agent 0.160 0.277 0.773 0.165 0.152 0.374
Sci-Reproducer 0.160 0.283 0.763 0.294 0.091 0.357
@@
We evaluate Sci-Reproducer on the SciReplicate-Bench  benchmark using 7 advanced LLMs, including five non-reasoning LLMs: GPT-4o-mini (4o mini, 2024), GPT-4o (GPT-4o, 2024), Claude-Sonnet-3.7 (Claude-Sonnet-3.7, 2025), Gemini-2.0-Flash (Gemini-2.0-Flash, 2024), and Deepseek-V3 (DeepSeek-AI et al., 2024), and different versions of the reasoning models O3-mini (o3 mini, 2024), i.e., three different levels of reasoning intensity. For the reasoning graph accuracy metric, node matching is performed using GPT-4o, which may introduce some randomness. To reduce this variability, we set the temperature to 0 and top-p to 1, ensuring more deterministic generation. The calculation is repeated three times, and we report the average score as the final result.